# Multistatic-Radar RCS-Signature Recognition of Aerial Vehicles: A Bayesian Fusion Approach

### v2

@@ Abstract
Index Terms:
@@
Acknowledgements
Research was sponsored by the Army Research Laboratory and was accomplished under Cooperative Agreement Number W911NF-23-2-0014. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Laboratory or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.
I INTRODUCTION
@@
UAVs are a class of aircraft that do not carry a human operator and fly autonomously or are piloted remotely [5]. Historically, UAVs were designed solely for military applications such as surveillance and reconnaissance, target acquisition, search and rescue, and force protection [6, 7]. Recently we have entered the ”Drone Age [8], aka personal glsuav area; where there has been a rapid increase in civilian UAV use for cinematography, tourism, commercial ads, real estate surveying and hot-spot / communications [9, 10].
@@ II RELATED WORK
III Multi-Static RCS Dataset Generation
@@
III-A Existing Data
We analyze the approximate RCS signatures of 7 different UAVs collected in [54]: the F450, Heli, Hexa, M100, P4P, Walkera, and Y600 . The data was collected in an anechoic chamber, where a quasi-monostatic radar transmitted EMWs at $F=15$ frequencies (26-40 GHz in 1 GHz steps) while stepped motors rotated the UAV around the azimuth axis ($\phi$ in Figure 1) and the elevation axis ($\theta$ in Figure 1) [54]. The azimuth spans from 0 to 180 degrees, while the elevation extends from -95 to 95 degrees, both in 1-degree increments. An example of the RCS image of a UAV from [54] is shown in Figure 2. We leverage the real RCS signatures to create training and testing datasets corresponding to simulated/synthetic radar locations and drone trajectories.
III-B Training Data Generation
We sample azimuth and elevation uniformly at random, where the spread is bounded by each drone’s minimum and maximum aspect angles from the experimental setup in [54]. The azimuth and elevation samples, $\phi$ and $\theta$ respectively, are mapped to a measured RCS signature $\sigma\in\mathcal{R}^{F}$ from [54], where we linearly interpolate the mapped RCS measurements for continuous azimuth and elevations not collected in [54]. For each UAV type $c\in\mathcal{C}_{train}=[1,\ldots,7]$, we generate $N_{train}$ samples such that $|\mathcal{D}_{train}|=|\mathcal{C}_{train}|N_{train}$. Training a discriminative ML classifier under the perspective of a single radar at a single time point enables efficient sampling and training. Generating training data for multistatic radar configurations would deal with the curse of dimensionality, where trajectories (a sample trajectory being multiple time points concatenated) compounded with multiple radars would exponentially increase the required number of samples to adequately cover the input space. In summary, the training dataset $\mathcal{D}_{train}$ has the following structure:
@@ III-B Training Data Generation
where $x_{i}^{(t)}=[\sigma_{i}^{(t)},\phi_{i}^{(t)},\theta_{i}^{(t)}]$ contains the noisy RCS signature, the noisy azimuth and the noisy elevation. The noisy RCS signature, azimuth, and elevation will be discussed in Section III.
@@
III-C Testing Data Generation
@@
where the positive $x$-axis of the target UAV coordinate frame is the UAV ”forward facing” direction, $e_{x}$ is the x-axis unit vector, and $\mathcal{U}$ is the UAV position. The initial position and yaw of the target UAV is sampled uniformly at random:
@@
where $\rho$ is the length of the RLOS to the target UAV. A single time step of a simulated UAV being tracked by multiple individual radars is depicted in Figure 3, featuring the displayed azimuths and elevations. The simulated azimuth and elevation time series data, $\phi^{(1:t)}\in\mathcal{R}^{L\times J}\text{ and }\theta^{(1:t)}\in\mathcal{R}^{L\times J}$ respectively, is mapped to the measured RCS time series data $z^{(1:t)}\in\mathcal{R}^{L\times(JF)}$. Due to the symmetry of the drones, the measured RCS for any azimuth not between $[0,180]$ is approximated as the RCS corresponding to $\phi=\phi+180$. We note a negative $\Delta z$ is used to align our target UAV coordinate frame with the experimental setup in [54], where the azimuth and elevation are calculated with respect to the bottom of the drone.
@@
We train the discriminative ML model on the RCS, azimuth, and elevation measurements described in Section III Subsection III-B, and evaluate our RATR framework on the trajectories described in Section III Subsection III-C. We next discuss how these discriminative ML models are integrated into the RATR framework.
IV METHODOLOGY
At each time step $t$ multiple radars pulse EMWs at the target UAV, where each radar $j$ calculates a RCS signature, azimuth, and elevation $x_{j}$. Each radar inputs its respective observation to a local discriminative ML model to output a UAV type probability vector. Here, local denontes that the ML resides at the individual radar hardware. All the individual radar UAV type probability vectors are combined by a fusion method (Section IV Subsection IV-B). Subsequently, the posterior UAV type probability distribution, based on the past RCS signatures $z^{(1:t-1)}$, is updated with the fused UAV type probability distribution from time step $t$. A RBC framework is applied such that the posterior probability distribution recursively updates as the radars continuously track a moving UAV. Our framework is shown in Figure 4.
Figure 4: RATR block diagram. Multiple radars continuously track and pulse an EMW at a moving UAV target. The each individual radar’s belief about the UAV type is at time step $t$ is fused in an Bayesian way and used to update recursively a posterior belief about the UAV type.
@@ IV METHODOLOGY
IV-A ML Models
@@ IV-B Fusion Methods
IV-B1 Bayesian Optimal Fusion
@@ IV-B5 Maximum
IV-C Recursive Bayesian Classification
@@ V EXPERIMENTS
Figure 5: One data sample from the training test data is shown for illustrative purposes. The first subfigure (from the left) depicts four uniformly spaced radars tracking and classifying a target UAV navigating with a random walk trajectory. The second subfigure illustrates the evolution of the RBC posterior probability for the UAV type over time as more RCS, azimuth, and elevation measurements are collected. The third and fourth figures display the change in elevation from ground level and the top view of the X-Y position of the UAV over time, respectively.
V-A Configuration
@@
V-B Signal-to-Noise Ratio
@@ V-C Results
Figure 6: Each figure axis, where each row specifies the SNR and each column specifies the ML model, plots the RBC accuracy over time of 16 radars (and a single radar) for the OBF, soft-voting, hard-voting, single radar, and maximum fusion methods. The azimuth and elevation jitter standard deviations was $\approx$23.
Figure 7: Each figure axis specifies the SNR, with the first row displaying results with azimuth and elevation having a jitter standard deviation of 23, and the second row displaying results without azimuth and elevation. The bar color represents different fusion methods, and the line atop the bar indicates the 95% confidence interval over 10 Monte Carlo trials. All results are for 16 radars in a multistatic configuration, except single radar.
@@
Figure 8: The RBC accuracy performance versus the number of radars increasing as a power of 2 (4, 16, 64 radars) for each discriminative ML model. The subfigures from left to the right increase the SNRdB and the term $J$ in the legend denotes the number of radars in a surveillance area.
@@
We visualize the RBC accuracy as the SNR changes (Figure 7). We observe an ”S” curve by following the top of the barplot across columnns of figure axes, where an increase in the signal power does not lead to an increase in classification performance at the upper extremes of SNR. However, radar channels typically operate in the 0 to -20 dB SNR range. When the RATR is operating at lower SNR ranges, having more observations of the target UAV substantially improves the classification performance versus a single radar (Figure 8).
Figure 9: Each subfigure is the RBC accuracy for 16 radars versus the azimuth and elevation standard deviation of the uniform distribution. The subfigures from left to the right decrease the SNRdB. Each line color denotes a different discriminative ML model, and the line marker denotes the fusion method.
If the SNR is high, the classification performance of a single radar should approach that of multi-radar classification, given a sufficiently large dwell time and suitable target geometries, especially when using a complex ML model (Figure 6).
@@
VI CONCLUSION
This paper is the first to introduce a fully Bayesian RATR for UAV type classification in multistatic radar configurations using RCS signatures. We evaluated the classification accuracy and robustness of our method across diverse SNR settings using RCS, azimuth, and elevation time series data generated by random walk drone trajectories. Our empirical results demonstrate that integrating the OBF method with RBC in multistatic radar signifcantly enhances ATR. Additionally, our method exhibits greater robustness in lower SNR, with larger relative improvements observed compared to single monostatic radar and other non-trivial fusion methods. Thus, fully Bayesian RATR in multistatic radar configurations using RCS signatures improves both classification accuracy and robustness.

### v3

# Student Member, IEEE

@@ Abstract
Code: https://github.com/mlpotter/RCS_ATR
{IEEEkeywords}
@@
1 INTRODUCTION
@@
UAVs are a class of aircraft that do not carry a human operator and fly autonomously or are piloted remotely [5]. Historically, UAVs were designed solely for military applications such as surveillance and reconnaissance, target acquisition, search and rescue, and force protection [6, 7]. Recently we have entered the “Drone Age[8], aka personal UAV area; where there has been a rapid increase in civilian UAV use for cinematography, tourism, commercial ads, real estate surveying and hot-spot / communications [9, 10].
@@ 2 RELATED WORK
3 Multi-Static RCS Dataset Generation
@@
3.1 Existing Data
We analyze the approximate RCS signatures of 7 different UAVs collected in [38]: the F450, Heli, Hexa, M100, P4P, Walkera, and Y600 . The data was collected in an anechoic chamber, where a quasi-monostatic radar transmitted EMWs at $F=15$ frequencies (26-40 GHz in 1 GHz steps) while stepped motors rotated the UAV around the azimuth axis ($\phi$ in Figure 2) and the elevation axis ($\theta$ in Figure 2) [38]. We note that UAVs were non-operational during the data collection process, and therefore the measured RCS signature of the drones in flight may slightly change due to the rotating propellers. The azimuth spans from 0 to 180 degrees, while the elevation extends from -95 to 95 degrees, both in 1-degree increments. An example of the RCS image of a UAV from [38] is shown in Figure 1. We leverage the real RCS signatures to create training and testing datasets corresponding to simulated/synthetic radar locations and drone trajectories.
3.2 Training Data Generation
We sample azimuth and elevation uniformly at random, where the spread is bounded by each drone’s minimum and maximum aspect angles from the experimental setup in [38]. The azimuth and elevation samples, $\phi$ and $\theta$ respectively, are mapped to a measured RCS signature $\sigma\in\mathcal{R}^{F}$ from [38], where we bilinearly interpolate the mapped RCS measurements for continuous azimuth and elevations not collected in [38]. For each UAV type $c\in\mathcal{C}_{train}=[1,\ldots,7]$, we generate $N_{train}$ samples such that $|\mathcal{D}_{train}|=|\mathcal{C}_{train}|N_{train}$. Training a discriminative ML classifier under the perspective of a single radar at a single time point enables efficient sampling and training. Generating training data for multistatic radar configurations would deal with the curse of dimensionality, where trajectories (a sample trajectory being multiple time points concatenated) compounded with multiple radars would exponentially increase the required number of samples to adequately cover the input space. In summary, the training dataset $\mathcal{D}_{train}$ has the following structure:
@@ 3.2 Training Data Generation
where $x_{i}^{(t)}=[\sigma_{i}^{(t)},\phi_{i}^{(t)},\theta_{i}^{(t)}]$ contains the noisy RCS signature, the noisy azimuth and the noisy elevation. The noisy RCS signature, azimuth, and elevation will be discussed in Section 3. Note that the sampling approach of bilinear interpolation and then addition of independent Gaussian noise ensures unique RCS measurements.
@@
Figure 4: The relationship between the target coordinate frame, the radar and RLOS, and the world coordinate frame. The x-axis, y-axis, and z-axis of the target UAV coordinate frame are the red, green, and blue arrows respectively with the x-axis as the forward direction. The position of the target is the translation with respect to the world coordinate frame.
3.3 Testing Data Generation
@@
where the positive $x$-axis of the target UAV coordinate frame is the UAV “forward facing” direction, $e_{x}$ is the x-axis unit vector, and $\mathcal{U}$ is the UAV position. Furthermore, $R_{\gamma},R_{\alpha},R_{\beta}$ are the Euler angle rotations for yaw, pitch, roll respectively of the target coordinate frame , and $d$ is the x-axis direction of the rotated target coordinate frame. The initial position and yaw of the target UAV is sampled uniformly at random:
@@
where $\rho$ is the length of the RLOS to the target UAV. Figure 4 visualizes the UAV coordinate frame highlighting the yaw, pitch, roll angles. The Figure 4 also shows how the radar line of sight intersects with the target coordinate frame to determine the azimuth and elevation angle. A single time step of a simulated UAV being tracked by multiple individual radars is depicted in Figure 3, featuring the displayed azimuths and elevations. The simulated azimuth and elevation time series data, $\phi^{(1:t)}\in\mathcal{R}^{L\times J}\text{ and }\theta^{(1:t)}\in\mathcal{R}^{L\times J}$ respectively, is mapped to the measured RCS time series data $z^{(1:t)}\in\mathcal{R}^{L\times(JF)}$. Due to the symmetry of the drones, the measured RCS for any azimuth not between $[0,180]$ is approximated as the RCS corresponding to $\phi=\phi+180$. We note a negative $\Delta z$ is used to align our target UAV coordinate frame with the experimental setup in [38], where the azimuth and elevation are calculated with respect to the bottom of the drone.
@@
We train the discriminative ML model on the RCS, azimuth, and elevation measurements described in Section 3 Subsection 3.2, and evaluate our RATR framework on the trajectories described in Section 3 Subsection 3.3. We next discuss how these discriminative ML models are integrated into the RATR framework.
4 METHODOLOGY
At each time step $t$ multiple radars pulse EMWs at the target UAV, where each radar $j$ calculates a RCS signature, azimuth, and elevation $x_{j}$. Each radar inputs its respective observation to a local discriminative ML model to output a UAV type probability vector. Here, local denotes that the ML resides at the individual radar hardware. All the individual radar UAV type probability vectors are combined by a fusion method (Section 4 Subsection 4.2). Subsequently, the posterior UAV type probability distribution, based on the past RCS signatures $z^{(1:t-1)}$, is updated with the fused UAV type probability distribution from time step $t$. A RBC framework is applied such that the posterior probability distribution recursively updates as the radars continuously track a moving UAV. Our framework is shown in Figure 5.
Figure 5: RATR block diagram. Multiple radars continuously track and pulse an EMW at a moving UAV target. The each individual radar’s belief about the UAV type is at time step $t$ is fused in an Bayesian way and used to update recursively a posterior belief about the UAV type. In the experiments section, we replace the red block (the OBF fusion method) with non-Bayesian fusion methods for comparison, noting that any probabilistic fusion method which produces a probability distribution over drone type can replace the red block.
@@ 4 METHODOLOGY
Figure 6: One sample (a single trajectory for the F450 drone) from the test dataset is shown for illustrative purposes. The first subfigure (from the left) depicts four uniformly spaced radars tracking and classifying a target UAV navigating with a random walk trajectory. The second subfigure illustrates the evolution of the RBC posterior probability for the UAV type over time as more RCS, azimuth, and elevation measurements are collected. The third and fourth figures display the change in elevation from ground level and the top view of the X-Y position of the UAV over time, respectively.
4.1 ML Models
@@ 4.2 Fusion Methods
4.2.1 Bayesian Optimal Fusion
@@ 4.2.5 Maximum
Figure 7: The RBC accuracy performance averaged across the test dataset versus the number of radars increasing as a power of 2 (4, 16, 64 radars) for each discriminative ML model. The subfigures from left to the right increase the SNRdB and the term $J$ in the legend denotes the number of radars in a surveillance area.
4.3 Recursive Bayesian Classification
@@ 5 EXPERIMENTS
5.1 Configuration
@@
    for experimental purposes, we assume that the targets are always detected and the RCS values are readily estimated (however, the methodology is independent of the detection procedure in radar pipelines).
  2. 2.
@@
Figure 8: Each figure axis specifies the SNR, with the first row displaying results averaged across the test dataset with azimuth and elevation having a jitter standard deviation of 23 degrees, and the second row displaying results averaged across the test dataset without azimuth and elevation input features. The bar color represents different fusion methods, and the line atop the bar indicates the 95% confidence interval over 10 Monte Carlo trials. All results are for 16 radars in a multistatic configuration, except single radar.
5.2 Signal-to-Noise Ratio
@@
Figure 9: Each figure axis, where each row specifies the SNR and each column specifies the ML model, plots the RBC accuracy averaged across the test dataset over time of 16 radars (and a single radar) for the OBF, soft-voting, hard-voting, single radar, and maximum fusion methods. The azimuth and elevation jitter standard deviations was $\approx$23 degrees.
@@ 5.3 Results
Figure 10: Each subfigure is the RBC accuracy averaged across the test dataset for 16 radars versus the azimuth and elevation standard deviation of the uniform distribution. The subfigures from left to the right decrease the SNRdB. Each line color denotes a different discriminative ML model, and the line marker denotes the fusion method.
@@
We visualize the RBC accuracy as the SNR changes (Figure 8). We observe an “S” curve by following the top of the barplot across columnns of figure axes, where an increase in the signal power does not lead to an increase in classification performance at the upper extremes of SNR. However, radar channels typically operate in the 0 to -20 dB SNR range. When the RATR is operating at lower SNR ranges, having more observations of the target UAV substantially improves the classification performance versus a single radar (Figure 7). If the SNR is high, the classification performance of a single radar should approach that of multi-radar classification, given a sufficiently large dwell time and suitable target geometries, especially when using a complex ML model (Figure 9).
@@
Although this paper focuses on the algorithmic perspective of UAV type classification, we note that without optimized code we achieve a 3.6 milliseconds (ms) inference time at a given time step on the CPU of a Zephyrus M16 laptop with a 12th Gen Intel(R) Core(TM) i9-12900H 2.5 GHz processor. That is, in 3.6 ms each local model outputs a probability distribution, the local model probability distributions are fused with the OBF method, and the running global posterior prediction is updated with RBC.
6 CONCLUSION
This paper is the first to introduce a fully Bayesian RATR for UAV type classification in multistatic radar configurations using RCS signatures. We evaluated the classification accuracy and robustness of our method across diverse SNR settings using RCS, azimuth, and elevation time series data generated by random walk drone trajectories. Our empirical results demonstrate that integrating the OBF method with RBC in multistatic radar significantly enhances ATR. Additionally, our method exhibits greater robustness in lower SNR, with larger relative improvements observed compared to single monostatic radar and other non-trivial fusion methods. Thus, fully Bayesian RATR in multistatic radar configurations using RCS signatures improves both classification accuracy and robustness.